body temps: mean of 96.20 standard deviation of .82
If you have a body temp of 97.00 degrees, what is your percentile score?

Answer by ewatrrr(24785)

 
Hi,
mean of 96.20 standard deviation of .82
P(X=97) = P(z=(97-96.2)/.82) = P(z= .9756) = .8354 or 83.54%
Excel function: NORMSDIST(0.9756)= .8354